differentiate the following equation using derivative rules:
f(x)= sqrt(x)
----------------------
2+x

Answer by Alan3354(69443)

differentiate the following equation using derivative rules:
f(x)= sqrt(x)
f(x) = x^(1/2)
f'(x) = (1/2)x^(-1/2)
= 1/(2sqrt(x))
= sqrt(x)/(2x)
----------------------
2+x
Seriously?
If you can't do this one, you're in the wrong class.
--------